OpenStudy (anonymous):

what is integral of sin wt^2 from 0 to T

OpenStudy (anonymous):

It's not an elementary function, unless you think the Fresnel S function is elementary.

OpenStudy (anonymous):

\[ S(T) = \int_0^T\sin(t^2)\;dt \]

OpenStudy (anonymous):

i want to know how it can be solved

OpenStudy (anonymous):

What do you mean by solved?

OpenStudy (anonymous):

Is it \(\sin (wt^2)\) or \(\sin^2(wt)\)

OpenStudy (ybarrap):

This might help - everything you need to evaluate it - http://functions.wolfram.com/GammaBetaErf/FresnelS/